The Average weight of 25 chocolate bars selected from a normally distributed population is 200 grams with a standard deviation of 10 grams. Find the point estimate and the confidence interval using 95%confidence level.

Steps: Solution:

2.Specify the confidence interval criteria.

b.Determine the test statistic to be used The test statistic is ___-test(4)

to calculate the interval.

c.State the level of confidence The level of confidence is _______(5)



1
Expert's answer
2022-04-28T16:06:21-0400

A sample mean is a point estimate of a population mean: "\bar{x}=200\ g."

A confidence interval is the most common type of interval estimate.

a. The critical value for "\alpha = 0.05" and "df = n-1 = 24" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.063899."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(200-2.063899\times\dfrac{10}{\sqrt{25}},"

"200+2.063899\times\dfrac{10}{\sqrt{25}})"

"=(195.8722, 204.1278)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "195.8722<\mu<204.1278," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(195.8722, 204.1278)."


b. The test statistic is t-test to calculate the interval.


c. The level of confidence is 0.95.